OpenStudy (anonymous):

Help me here Please: Step-By-Step Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5). Do not use spaces and enter fractions as improper reduced fractions when needed.

OpenStudy (anonymous):

@waterineyes @Calcmathlete @lgbasallote @mathslover or anyone?

OpenStudy (helder_edwin):

you have an equation of the form \[ y=mx+c \] so m=3

OpenStudy (anonymous):

@helder_edwin I already know the formula..

OpenStudy (anonymous):

Parallel Lines have Same Slope I think..

OpenStudy (helder_edwin):

two parallel lines have different or same slope?

OpenStudy (anonymous):

So, m= 3.. \[y - 2.5 = 3(x + 10)\] Solve for y...

OpenStudy (lgbasallote):

@Trexy can you clarify what you need help on..

OpenStudy (anonymous):

@lgbasallote Kuya hindi ko po alam kung paano po ito sasagutin kaya tinanung ko kayong lahat.. Sorry mahina po talaga ako sa Math..

OpenStudy (lgbasallote):

you said you already know the formula...so what dont you know?

OpenStudy (anonymous):

NIce pic @lgbasallote

OpenStudy (anonymous):

@Trexy I have written the equation above can you solve for y there??

OpenStudy (lgbasallote):

hah you too

OpenStudy (lgbasallote):

lol

OpenStudy (anonymous):

thank you..

OpenStudy (anonymous):

Ha ha ha..

OpenStudy (anonymous):

I am asking you and not demanding to say thank you to me @Trexy

OpenStudy (anonymous):

hahah.. Jeez.. You, Zepp, and Cal are Scary..